### v1

@@ Abstract
We investigate the semigroup of integer points inside a convex cone. We extend classical results in integer linear programming to integer conic programming. We show that the semigroup associated with nonpolyhedral cones can sometimes have a notion of finite generating set. We show this is true for the cone of positive semidefinite matrices (PSD) and the second-order cone (SOC). Both cones have a finite generating set of integer points, similar in spirit to Hilbert bases, under the action of a finitely generated group. We also extend notions of total dual integrality, Gomory-Chvátal closure, and Carathéodory rank to integer points in arbitrary cones.
Keywords:
Integer Points Convex Cones Semigroups Hilbert bases Conic Programming Positive Semidefinite Cone Second-Order Cone
@@ 1 Introduction
A semigroup $S$ is a subset of $\mathbb{Z}^{N}$ that contains $\mathbf{0}$ and is closed under addition. Given a convex cone $C\subseteq\mathbb{R}^{N}$, the integer points $S_{C}\mathrel{\mathop{\ordinarycolon}}=C\cap\mathbb{Z}^{N}$ form a semigroup which we will call the conical semigroup of $C$. In particular, given any compact convex body $K\subseteq\mathbb{R}^{n}$, the integer points $\text{cone}(K\times\{1\})\cap\mathbb{Z}^{n+1}$ form a conical semigroup. Conical semigroups appear not just in optimization [1, 6], but also in algebra and number theory [2, 3]. Given a convex cone $C\subseteq\mathbb{R}^{N}$ for $N\geq 1$, we say a subset $B\subseteq S_{C}$ is an integral generating set of $S_{C}$ if for any $s\in S_{C}$ there exist $b_{1},\dots,b_{m}\in B$ and $c_{1},\dots,c_{m}\in\mathbb{Z}_{\geq 0}$ such that $s=\sum_{i=1}^{m}c_{i}b_{i}$, for some $m\geq 1$. Furthermore, we call $B$ a conical Hilbert basis if $B$ is an inclusion-minimal integral generating set.
When the defining cone $C$ is rational, polyhedral and pointed, there is abundant literature on the topic. It is well-known that we have a unique finite Hilbert basis in this case [11, 20]. Historically, Hilbert bases have been fundamental in the theory and algorithms of combinatorial optimization. For example, determining if a rational system $A\mathbf{x}\leq\mathbf{b}$ is totally dual integral (TDI) is equivalent to checking if, for every face $F$ of the polyhedron $P\mathrel{\mathop{\ordinarycolon}}=\{\mathbf{x}\mathrel{\mathop{\ordinarycolon}}A\mathbf{x}\leq\mathbf{b}\}$, the rows of $A$ which are active in a face $F$ form a Hilbert basis for cone($F$)[20].
It is natural to ask, what properties transfer from rational polyhedral cones to arbitrary convex cones? For instance, do we preserve any notion of finiteness in generating sets for semigroups when we relax the polyhedral condition and instead consider general conical semigroups? Are there Hilbert bases for general cones? This paper discusses finite generation for conical semigroups and extends the polyhedral cone theory of Hilbert bases to nonpolyhedral convex cones. Our main results will pertain to the semigroups arising from the cone of positive semidefinite matrices and the second-order cone. Both cones play a key role in modern optimization [4, 5]. We also discuss some applications of our nonpolyhedral point of view.
In what follows, we denote $\mathrm{GL}(N,\mathbb{Z})\mathrel{\mathop{\ordinarycolon}}=\{U\in\mathbb{Z}^{N\times N}\mathrel{\mathop{\ordinarycolon}}~{}|\det(U)|=1\}$. Here is a new notion of finite generation for conical semigroups.
Definition 1
@@
Note that when $C$ is a (pointed) rational polyhedral cone, then the conical semigroup $S_{C}=C\cap\mathbb{Z}^{N}$ is $(R,G)$-finitely generated by $R$, its Hilbert basis, and $G$, the trivial group $\{I_{N}\}$. Similarly, note that if $S_{C}$ is an $(R,G)$-finitely generated semigroup, then $\cup_{r\in R}G\cdot r$ is an integral generating set of $S_{C}$, which is a superset of a conical Hilbert basis. We call $R$ the set of roots of $S_{C}$, and $\cup_{r\in R}G\cdot r$ the set of generators for $S_{C}$.
While a nonpolyhedral cone cannot be finitely generated in the usual sense, using a possibly infinite (finitely generated) group $G$ allows us to extend our understanding beyond the polyhedral case. Because the possibly infinite generators for $S_{C}$ can be obtained by group action $G$ on a finite set $R$ and $G$ is finitely generated, this allows for the possibility of algorithmic methods. The well-known Krein-Milman theorem states that any point in a closed pointed cone $C$ can be generated by extreme rays, denoted by $\operatorname{ext}(C)$ [4]. When we restrict to the conical semigroup $S_{C}$ and nonnegative integer combinations, the primitive integer points on the extreme rays of $C$ must be contained in the set of generators of $S_{C}$, where an integer point $x=(x_{1},\dots,x_{N})\in\mathbb{Z}^{N}$ is primitive if $\mathrm{gcd}(x_{1},\dots,x_{N})=1$. We call the integer points of $S_{C}$ on the extreme rays of $C$ extreme points, denoted by $\operatorname{ext}(S_{C})\mathrel{\mathop{\ordinarycolon}}=\{y\in S_{C}\mathrel{\mathop{\ordinarycolon}}y\in\operatorname{ext}(C)\cap\mathbb{Z}^{N}\}$. However, as in the polyhedral case, the generators will often include extra nonextreme boundary points or even interior points. We provide the following definition of sporadic points that cannot have an extreme point subtracted from them and still remain within the cone.
Definition 2
@@
The two convex cones of interest in this work are positive semidefinite cone (PSD) and second-order cone (SOC). In Sections 2 and 3 of this paper, we will present the following two main results pertaining to integer points in the PSD cone $\mathcal{S}^{n}_{+}(\mathbb{Z})$, and those in the SOC $\mathrm{SOC}(n)\cap\mathbb{Z}^{n}$.
Theorem 1.1
@@
Theorem 1.2
For dimension $3\leq n\leq 10$, the conical semigroup $\mathrm{SOC}(n)\cap\mathbb{Z}^{n}$ is $(R,G)$-finitely generated with $G$ and $R$ defined in Section 3.
We say that two matrices $X_{1},X_{2}$ are unimodularly equivalent if $X_{2}=U\cdot X_{1}=UX_{1}U^{\mathsf{T}}$ for some $U\in\mathrm{GL}(n,\mathbb{Z})$. It is easy to see that it defines an equivalence relation for all integer PSD matrices. Note that the equivalence class of $\mathbf{e}_{1}\mathbf{e}_{1}^{\mathsf{T}}$ are all rank-1 integer matrix $\mathbf{x}\mathbf{x}^{\mathsf{T}}$ for some primitive integer vector $\mathbf{x}\in\mathbb{Z}^{n}$. An interpretation of Theorem 1.1 is that for dimension $n\leq 5$, every integer PSD matrix can be represented as the sum of rank-$1$ matrices $\mathbf{x}\mathbf{x}^{\mathsf{T}}$ for some primitive integer vector $\mathbf{x}\in\mathbb{Z}^{n}$. However, the same result fails for dimension $n=6$. In this case, we will have that every integer PSD matrix can be represented as the sum of rank-$1$ matrices and one sporadic matrix $Y$, which is unimodularly equivalent to $M$ (this matrix was first found by [19]). In general, every integer PSD matrix can be represented as the sum of rank-$1$ matrices and one sporadic matrix, which is unimodularly equivalent to a matrix in the finite set $R$. Regarding prior work that inspired us, we mention [16] that contains a similar rank-1 decomposition structure for PSD $\{0,1\}$ matrices: a PSD $\{0,1\}$ matrix $X\in\mathcal{S}^{n}_{+}(\mathbb{Z})\cap\{0,1\}^{n\times n}$ satisfies $X=\sum_{i\in K}\mathbf{x}_{i}\mathbf{x}_{i}^{\mathsf{T}}$ for $\mathbf{x}_{i}\in\{0,1\}^{n}$, where $K$ is a finite index set. Similarly, [18] extended the results to PSD $\{0,\pm 1\}$ matrices: a PSD $\{0,\pm 1\}$ matrix $X\in\mathcal{S}^{n}_{+}(\mathbb{Z})\cap\{0,\pm 1\}^{n\times n}$ satisfies $X=\sum_{i\in K}\mathbf{x}_{i}\mathbf{x}_{i}^{\mathsf{T}}$ for $\mathbf{x}_{i}\in\{0,\pm 1\}^{n}$, where $K$ is a finite index set. Our results extend to all integer positive semidefinite matrices. For the second-order cone, we extended the construction of the Barning-Hall tree in [8] for the primitive extreme points (or Pythagorean tuples) to classify the sporadic points.
While it might be tempting to believe that these results hint that all conical semigroups are $(R,G)$-finitely generated for some finite set $R$ and some subgroup $G\subseteq\mathrm{GL}(N,\mathbb{Z})$, we conjecture the contrary:
Conjecture 1
There exists a conical semigroup $S_{C}$ that is not $(R,G)$-finitely generated for any choice of $R$ and $G$.
@@
The first application regards the notion of Chvátal-Gomory cuts which is useful in the branch-and-cut methods for integer programming. How much of this can be extended to conic integer programming? Given a linear map $\mathcal{A}\mathrel{\mathop{\ordinarycolon}}\mathbb{R}^{m}\to\mathbb{R}^{N}$ and $\mathbf{c}\in\mathbb{R}^{N}$, we define a linear conical inequality (LCI) system as
$\mathrm{LCI}_{C}(\mathbf{c},\mathcal{A})\mathrel{\mathop{\ordinarycolon}}=\{\mathbf{x}\in\mathbb{R}^{m}\mathrel{\mathop{\ordinarycolon}}\mathbf{c}-\mathcal{A}(\mathbf{x})\in C\}$
@@
To approximate the convex hull of $Z\mathrel{\mathop{\ordinarycolon}}=\mathrm{LCI}_{C}(\mathbf{c},\mathcal{A})\cap\mathbb{Z}^{m}$, a commonly used approach (quite similar to its polyhedral version) is to add Chvátal-Gomory (CG) cuts, which are defined as follows [17]. If $\mathbf{u}\in\mathbb{Z}^{m}$ is an integral vector and $v\in\mathbb{R}$ a real number such that the linear inequality $\mathbf{u}^{\mathsf{T}}\mathbf{x}\leq v$ is valid for all $x\in\mathrm{LCI}_{C}(c,\mathcal{A})$, then the inequality $\mathbf{u}^{\mathsf{T}}\mathbf{x}\leq\lfloor v\rfloor$ is valid for all $\mathbf{x}\in Z$ and called a CG cut. There are possibly infinitely many CG cuts so we define the (elementary) CG closure as
$\mathrm{CG}\text{-}\mathrm{cl}(Z)\mathrel{\mathop{\ordinarycolon}}=\bigcap_{\begin{subarray}{c}(\mathbf{u},v)\in\mathbb{Z}^{m}\times\mathbb{R}\mathrel{\mathop{\ordinarycolon}}\\ S\subseteq\{\mathbf{x}\mathrel{\mathop{\ordinarycolon}}\mathbf{u}^{\mathsf{T}}\mathbf{x}\leq v\}\end{subarray}}\left\{\mathbf{x}\in\mathbb{R}^{m}\mathrel{\mathop{\ordinarycolon}}\mathbf{u}^{\mathsf{T}}\mathbf{x}\leq\lfloor v\rfloor\right\}.$ (1)
@@
as, by definition, $w\circ\mathcal{A}(\mathbb{Z}^{m})\in\mathbb{Z}$. Conversely, if the conical semigroup $S_{C^{*}}\mathrel{\mathop{\ordinarycolon}}=C^{*}\cap\mathbb{Z}^{N}$ is $(R,G)$-finitely generated, then we can get all CG cuts through $R$ and $G$ for our TDI LCI system. This is one of the nice consequences of this property.
Theorem 1.3
Suppose $C\subset\mathbb{R}^{N}$ is a full-dimensional convex cone such that $S_{C^{*}}\mathrel{\mathop{\ordinarycolon}}=C^{*}\cap\mathbb{Z}^{N}$ is $(R,G)$-finitely generated, and $\mathrm{LCI}_{C}(\mathbf{c},\mathcal{A})$ is TDI. Then the CG closure for $Z\mathrel{\mathop{\ordinarycolon}}=\mathrm{LCI}_{C}(\mathbf{c},\mathcal{A})\cap\mathbb{Z}^{m}$ can be described by
$\mathrm{CG}\text{-}\mathrm{cl}(Z)=\left\{\mathbf{x}\in\mathbb{R}^{m}\mathrel{\mathop{\ordinarycolon}}(g\cdot r)^{\mathsf{T}}\mathcal{A}(\mathbf{x})\leq\lfloor(g\cdot r)^{\mathsf{T}}\mathbf{c}\rfloor,\quad\forall\,r\in R,g\in G\right\}.$
@@
Theorem 1.4
Let $C\subset\mathbb{R}^{N}$ be an arbitrary pointed convex cone and $S_{C}\mathrel{\mathop{\ordinarycolon}}=C\cap\mathbb{Z}^{N}$. Then $\mathrm{ICR}(S_{C})\leq 2N-2$.

### v2

@@ Abstract
We investigate the semigroup of integer points inside a convex cone. We extend classical results in integer linear programming to integer conic programming. We show that the semigroup associated with nonpolyhedral cones can sometimes have a notion of finite generating set with the help of a group action. We show this is true for the cone of positive semidefinite matrices (PSD) and the second-order cone (SOC). Both cones have a finite generating set of integer points, similar in spirit to Hilbert bases, under the action of a finitely generated group. We also extend notions of total dual integrality, Gomory-Chvátal closure, and Carathéodory rank to integer points in arbitrary cones.  [^2] This research is partially based upon work supported by the National Science Foundation under Grant No. DMS-1929284 while the first, third, and fourth authors were in residence at the Institute for Computational and Experimental Research in Mathematics in Providence, RI, during the Discrete Optimization program. The first and second authors were partially supported by the National Science Foundation under grants No. DMS 1818969 and No. DMS 2348578. We thank Kurt Anstreicher, Greg Blekherman, Santanu Dey, Renata Sotirov, Pablo Parrilo, Chiara Meroni, and Bento Natura for relevant comments.
Keywords: Integer Points, Convex Cones, Semigroups, Hilbert bases, Conic Programming, Positive Semidefinite Cone, Second-Order Cone
volume: (volume: (volume: (
@@ 1 Introduction
A semigroup $S$ is a subset of $\mathbb{Z}^{N}$ that contains $\mathbf{0}$ and is closed under addition. Given a convex cone $C\subseteq\mathbb{R}^{N}$, the integer points $S_{C}:=C\cap\mathbb{Z}^{N}$ form a semigroup which we will call the conical semigroup of $C$. In particular, given any compact convex body $K\subseteq\mathbb{R}^{n}$, the integer points $\text{cone}(K\times\{1\})\cap\mathbb{Z}^{n+1}$ form a conical semigroup. Conical semigroups appear not just in optimization [2, 3], but also in algebra and number theory [4, 5]. Given a convex cone $C\subseteq\mathbb{R}^{N}$ for $N\geq 1$, we say a subset $B\subseteq S_{C}$ is an integral generating set of $S_{C}$ if for any $s\in S_{C}$ there exist $b_{1},\dots,b_{m}\in B$ and $c_{1},\dots,c_{m}\in\mathbb{Z}_{\geq 0}$ such that $s=\sum_{i=1}^{m}c_{i}b_{i}$, for some $m\geq 1$. Furthermore, we call $B$ a conical Hilbert basis if $B$ is an inclusion-minimal integral generating set.
When the defining cone $C$ is rational, polyhedral, and pointed, there is abundant literature on the topic. It is well-known that we have a unique finite Hilbert basis in this case [6, 7]. Historically, Hilbert bases have been fundamental in the theory and algorithms of combinatorial optimization. For example, determining if a rational system $A\mathbf{x}\leq\mathbf{b}$ is totally dual integral (TDI) is equivalent to checking if, for every face $F$ of the polyhedron $P:=\{\mathbf{x}:A\mathbf{x}\leq\mathbf{b}\}$, the rows of $A$ which are active in a face $F$ form a Hilbert basis for cone($F$)[7].
It is natural to ask, what properties transfer from rational polyhedral cones to arbitrary convex cones? Unlike for polyhedra, prior work on understanding the lattice points of general convex cones is very sparse. For instance, when is the convex hull of the lattice points a closed set? This has been investigated in [8]. The question we will consider is do we preserve any notion of finiteness in generating sets for semigroups when we relax the polyhedral condition and instead consider general conical semigroups? Are there Hilbert bases for general cones? In [9] the authors presented conditions for when a set of lattice points have a finite integral generating set, partially emulating polyhedra. Unfortunately, this is rarely the case. Our paper discusses finite generation for conical semigroups and extends the polyhedral cone theory of Hilbert bases to nonpolyhedral convex cones. Our first contribution to the subject is to add a group action, allowing for finite generation with the help of the group. Our main results will show that with a group action, the semigroups of the cone of positive semidefinite matrices and the second-order cone are finitely generated. Both cones play a key role in modern optimization [10, 11]. We also discuss some applications of our nonpolyhedral point of view.
In what follows, we denote $\mathrm{GL}(N,\mathbb{Z}):=\{U\in\mathbb{Z}^{N\times N}:~{}|\det(U)|=1\}$. Here is our new notion of finite generation for conical semigroups.
Definition 1.
@@
When $C$ is a (pointed) rational polyhedral cone, then the conical semigroup $S_{C}=C\cap\mathbb{Z}^{N}$ is $(R,G)$-finitely generated by $R$, its Hilbert basis, and $G$, the trivial group $\{I_{N}\}$. Similarly, note that if $S_{C}$ is an $(R,G)$-finitely generated semigroup, then $\cup_{r\in R}G\cdot r$ is an integral generating set of $S_{C}$, which is a superset of a conical Hilbert basis. We call $R$ the set of roots of $S_{C}$, and $\cup_{r\in R}G\cdot r$ the set of generators for $S_{C}$.
While a nonpolyhedral cone cannot be finitely generated in the usual sense, using a possibly infinite (finitely generated) group $G$ allows us to extend our understanding beyond the polyhedral case. Because the possibly infinite generators for $S_{C}$ can be obtained by group action $G$ on a finite set $R$ and $G$ is finitely generated, this allows for the possibility of algorithmic methods. The well-known Krein-Milman theorem states that any point in a closed pointed cone $C$ can be generated by extreme rays, denoted by $\operatorname{ext}(C)$ [10]. When we restrict to the conical semigroup $S_{C}$ and nonnegative integer combinations, the primitive integer points on the extreme rays of $C$ must be contained in the set of generators of $S_{C}$, where an integer point $x=(x_{1},\dots,x_{N})\in\mathbb{Z}^{N}$ is primitive if $\mathrm{gcd}(x_{1},\dots,x_{N})=1$. We call the integer points of $S_{C}$ on the extreme rays of $C$ extreme points, denoted by $\operatorname{ext}(S_{C}):=\{y\in S_{C}:y\in\operatorname{ext}(C)\cap\mathbb{Z}^{N}\}$. However, as in the polyhedral case, the generators will often include extra nonextreme boundary points or even interior points. We provide the following definition of sporadic points which are points that cannot have an extreme point subtracted from them and still remain within the cone.
Definition 2.
@@
The two convex cones of interest in this work are the positive semidefinite cone (PSD) and the second-order cone (SOC). In Sections 2 and 3 of this paper, we will present the following two main results pertaining to integer points in the PSD cone, denoted $\mathcal{S}^{n}_{+}(\mathbb{Z})$, and the integer points in the second order cone, denoted $\mathrm{SOC}(n)\cap\mathbb{Z}^{n}$.
Theorem 1.
@@
We say that two matrices $X_{1},X_{2}$ are unimodularly equivalent if $X_{2}=U\cdot X_{1}=UX_{1}U^{\mathsf{T}}$ for some $U\in\mathrm{GL}(n,\mathbb{Z})$. It is easy to see that it defines an equivalence relation for all integer PSD matrices. Note that the equivalence class of $\mathbf{e}_{1}\mathbf{e}_{1}^{\mathsf{T}}$ are all rank-1 integer matrix $\mathbf{x}\mathbf{x}^{\mathsf{T}}$ for some primitive integer vector $\mathbf{x}\in\mathbb{Z}^{n}$. An interpretation of Theorem 1 is that for dimension $n\leq 5$, every integer PSD matrix can be represented as the sum of rank-$1$ matrices $\mathbf{x}\mathbf{x}^{\mathsf{T}}$ for some primitive integer vector $\mathbf{x}\in\mathbb{Z}^{n}$. However, the same result fails for dimension $n=6$. In this case, we will have that every integer PSD matrix can be represented as the sum of rank-$1$ matrices and one sporadic matrix $Y$, which is unimodularly equivalent to $M$ (this matrix was first found by [12]). In general, every integer PSD matrix can be represented as the sum of rank-$1$ matrices and one sporadic matrix, which is unimodularly equivalent to a matrix in the finite set $R$. Regarding prior work that inspired us, we mention [13] that contains a similar rank-1 decomposition structure for PSD $\{0,1\}$ matrices: a PSD $\{0,1\}$ matrix $X\in\mathcal{S}^{n}_{+}(\mathbb{Z})\cap\{0,1\}^{n\times n}$ satisfies $X=\sum_{i\in K}\mathbf{x}_{i}\mathbf{x}_{i}^{\mathsf{T}}$ for $\mathbf{x}_{i}\in\{0,1\}^{n}$, where $K$ is a finite index set. Similarly, [14] extended the results to PSD $\{0,\pm 1\}$ matrices: a PSD $\{0,\pm 1\}$ matrix $X\in\mathcal{S}^{n}_{+}(\mathbb{Z})\cap\{0,\pm 1\}^{n\times n}$ satisfies $X=\sum_{i\in K}\mathbf{x}_{i}\mathbf{x}_{i}^{\mathsf{T}}$ for $\mathbf{x}_{i}\in\{0,\pm 1\}^{n}$, where $K$ is a finite index set. Our results extend to all integer positive semidefinite matrices.
For the second-order cone family we were able to prove a similar theorem for dimensions up to $n=10$ by extending the construction of the Barning-Hall tree in [15] for the primitive extreme points (or Pythagorean tuples) to classify the sporadic points.
Theorem 2.
For dimension $3\leq n\leq 10$, the conical semigroup $\mathrm{SOC}(n)\cap\mathbb{Z}^{n}$ is $(R,G)$-finitely generated. The group $G$ and the roots $R$ will be defined in Section 3.
While it might be tempting to believe that all conical semigroups are $(R,G)$-finitely generated for some finite set $R$ and some subgroup $G\subseteq\mathrm{GL}(N,\mathbb{Z})$, this is not true in general. In Section 4 we present Example 1 which is a non-rational cone in the plane such that its semigroup is not $(R,G)$-finitely generated. For future work we leave open the following natural questions: which convex cones are (R,G)-finitely generated? Can we characterize them, at least for the plane?
@@
The first application regards the notion of Chvátal-Gomory cuts which is useful in the branch-and-cut methods for integer programming. How much of this can be extended to conic integer programming? Given a linear map $\mathcal{A}:\mathbb{R}^{m}\to\mathbb{R}^{N}$ and $\mathbf{c}\in\mathbb{R}^{N}$, we define a linear conical inequality (LCI) system as
$\mathrm{LCI}_{C}(\mathbf{c},\mathcal{A}):=\{\mathbf{x}\in\mathbb{R}^{m}:\mathbf{c}-\mathcal{A}(\mathbf{x})\in C\}$
@@
To approximate the convex hull of $Z:=\mathrm{LCI}_{C}(\mathbf{c},\mathcal{A})\cap\mathbb{Z}^{m}$, a commonly used approach (quite similar to its polyhedral version) is to add Chvátal-Gomory (CG) cuts, which are defined as follows [20]. If $\mathbf{u}\in\mathbb{Z}^{m}$ is an integral vector and $v\in\mathbb{R}$ a real number such that the linear inequality $\mathbf{u}^{\mathsf{T}}\mathbf{x}\leq v$ is valid for all $x\in\mathrm{LCI}_{C}(c,\mathcal{A})$, then the inequality $\mathbf{u}^{\mathsf{T}}\mathbf{x}\leq\lfloor v\rfloor$ is valid for all $\mathbf{x}\in Z$ and called a CG cut. There are possibly infinitely many CG cuts so we define the (elementary) CG closure as
$\mathrm{CG}\text{-}\mathrm{cl}(Z):=\bigcap_{\begin{subarray}{c}(\mathbf{u},v)\in\mathbb{Z}^{m}\times\mathbb{R}:\\ S\subseteq\{\mathbf{x}:\mathbf{u}^{\mathsf{T}}\mathbf{x}\leq v\}\end{subarray}}\left\{\mathbf{x}\in\mathbb{R}^{m}:\mathbf{u}^{\mathsf{T}}\mathbf{x}\leq\lfloor v\rfloor\right\}.$ (1)
@@
as, by definition, $w\circ\mathcal{A}(\mathbb{Z}^{m})\in\mathbb{Z}$. Conversely, if the conical semigroup $S_{C^{*}}:=C^{*}\cap\mathbb{Z}^{N}$ is $(R,G)$-finitely generated, then we can get all CG cuts through $R$ and $G$ for our TDI LCI system. This is one of the nice consequences of this property.
Theorem 3.
Suppose $C\subset\mathbb{R}^{N}$ is a full-dimensional convex cone such that $S_{C^{*}}:=C^{*}\cap\mathbb{Z}^{N}$ is $(R,G)$-finitely generated, and $\mathrm{LCI}_{C}(\mathbf{c},\mathcal{A})$ is TDI. Then the CG closure for $Z:=\mathrm{LCI}_{C}(\mathbf{c},\mathcal{A})\cap\mathbb{Z}^{m}$ can be described by
$\mathrm{CG}\text{-}\mathrm{cl}(Z)=\left\{\mathbf{x}\in\mathbb{R}^{m}:(g\cdot r)^{\mathsf{T}}\mathcal{A}(\mathbf{x})\leq\lfloor(g\cdot r)^{\mathsf{T}}\mathbf{c}\rfloor,\quad\forall\,r\in R,g\in G\right\}.$
@@
Theorem 4.
Let $C\subset\mathbb{R}^{N}$ be an arbitrary pointed convex cone and $S_{C}:=C\cap\mathbb{Z}^{N}$. Then $\mathrm{ICR}(S_{C})\leq 2N-2$.